ROTATIONAL SPECTRA AND STRUCTURE OF PHENYLACETYLENE-WATER COMPLEX

MAUSUMI GOSWAMI, L. NARASIMHAN, S. T. MANJU and \underline{E. ARUNAN}, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560012, India.

Rotational Spectra of phenylacetylene-water complex and its HDO and D_{2}O isotopomers have been measured. The spectra resemble an asymmetric top with κ = −0.73 for the parent isotopomer. Both ‘a’ and ‘b’ dipole transitions have been observed. All the transitions are split into two resulting from a possible internal rotation of H_{2}O in the complex. The measured rotational constants are A = 2672.0931(31)MHz, B = 996.3581(8)MHz, C = 731.7056(4)MHz for the stronger series and A = 2673.1331(44)MHz, B = 996.3926(10)MHz, C = 731.5737(6)MHz for the weaker series. Spectra for HDO isotopomer agree with a structure where water is in the plane of phenylacetylene by donating one of its hydrogen to the acetylene π cloud and the oxygen is involved in a secondary interaction forming C-H—O hydrogen bond with the ring hydrogen ortho to the acetylenic group. The experiments for D_{2}O isotopomer are in progress. The structure agrees with the one proposed from Fluorescence Dip Infrared Spectroscopy studies [1].

1. G. N. Patwari and co-workers, Private Communication.